A Complete Bibliography of Publications in The Rutherford Journal

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 July 2019 Version 1.01

Title word cross-reference

1951 [Car18].

After [Tee07]. Again [Sma18]. Agent [He06]. Alan [CP12, Cop18]. Analogue [Car07]. Anticipatory [Bod07]. Argument [Mus07]. Artificial [Rou18]. Atomic [Cha06]. Australian [BC07]. Automatic [Dor07, Swa10].

Babbage [Swa10]. Back [Sma18]. be [Sce18]. becomes [Mag06]. Behavior [Dre10]. Bletchley [Cop10, Wei06]. Bombe [Car10]. Breaking [Cop10]. Brief [Bow10, Sce18]. Business [Asp07].

C [BSS18]. Canterbury [Cla06]. Causality [He06]. Charles [Swa10]. Chemistry [Cha06, He06]. Chronology [Car07]. Civilization [Bon18]. Code [Cop10]. cognition [Ste03]. College [Cla06]. Colossus [Cop10]. Common [McC10]. Computation [Dre10, Mah10, Swa10]. Computational [Swa10]. Computer [CP12, Sha10]. Computers [Cor10, DT18]. Computing [BC07, Car07, CPSB10, T^{+}06]. Conference [Car18]. Contributors [Ano18b]. Creation [Bro06a]. Critical [Bro06a, Bro07].

D [Bro07]. Dalton [Cha06]. Darwinian [Bro06a]. Depend [He06]. designed [Hau06]. Dialogue [CC07]. Discovery [ADP06]. Do [Cas07, He06].

Early [Bow10]. Eggs [Wei06]. Elastician [Mag06]. Electronic [Cor10]. Else [Cas07]. Enabled [Asp10]. Events [Cas07]. Evolution [Bro06a, Ste03]. Existence [Ste18]. Experimental [Cha06]. Explanation [He06]. Export [Asp07].

Father [CP12]. Figure [Bon18]. First [Dor07]. Fish [Wei06]. Formalizing [McC10]. Formats [He06]. friendly [CK06]. Fujita [Asp07].

German [Cop10]. Grey [Bod07].

Hauser [Bro07]. Henry [Dor18]. Heritage [DT18]. History [BC07, Bow10, Cop10, CPSB10]. Hodsdon [Dor18]. Hollerith [ADP06]. Hostile [Bro06a, Ste03]. Human [Bro06a, Cor10, Ste03]. Hunting [Cor10].

Illustrated [Cop10]. Indian [Asp10]. Industry [Asp10]. Information [Bro06b]. Instinct [Bro06b]. Intelligence [Rou18]. Interpretation [Pro06]. Israel [Sha10]. Issue [CPSB10]. IT-Enabled [Asp10].

J [BSS18]. J. [BSS18]. Japanese [Asp07]. Jevons [BC06]. Journal [Ano18a].

Karl [Mag06]. Kim [Bro06a]. Knowledge [McC10]. Konrad [Bro06b].

Laid [Wei06]. Limitations [Cha06]. Logic [BC06, McC10, Num10]. look [Sce18]. Lorenz [Bro06b]. Lost [Cop18].

Machine [ADP06, Num10, T^{+}06]. Machines [Rou18]. Made [DT18]. Makudonaldo [Asp07]. Map [Tee07]. ’Marble [Dor18]. Marc [Bro07]. Mark [Car18]. Mathematical [Mah10, McC10]. Mathematics [CC07]. Meeting [Car18]. Memory [Dre10]. Metaphysics [Sma18]. Method [Swa10]. Mind [Bro06a, T^{+}06]. minds [Hau06, Bro07]. Miracle [Mus07]. Missed [Car18]. Modern [CP12, Mag06]. Moral [Hau06, Bro07]. Morality [Bro07]. Morphogenesis [Ric06]. Museums [Bow10]. Myth [Bro06a].

Named [Tee07]. Nature [Mah10, Hau06]. No [Wei06]. Notebook [Cop18]. Notice [Bro06a, Bro07]. Numbers [Cor10].

Obituary [BSS18]. Online [Bow10]. Organization [Num10]. Origins [Bon18, Mag06]. Our [DT18, Hau06].

P [Rou18]. P-Type [Rou18]. Paris [ADP06, Car18]. Park [Cop10, Wei06]. Particle [Yoc18]. Pearson [Mag06]. Periodic [Sce18]. Perspective [Yoc18]. Phylogenetic [Bro06b]. Physics [CC07, Sma18, Yoc18]. Piano [BC06]. Places [Tee07]. Policy [Asp10]. Preserving [DT18]. Prime [Cor10]. Prometheus [Bon18]. proposals [Sce18]. Prototype [ADP06]. published [Sce18].

Radio [Asp07]. Realism [Mus07]. Regarded [Sce18]. Reign [BC07]. Richards [Ric06]. right [Hau06]. Rise [Asp10]. Role [Asp10]. Rutherford [Cla06, Ano18a].

Science [Bro06a, Bro07, Cha06, Sha10, Tee07]. Scientific [Mus07]. Scientists [Tee07]. See [Cas07]. Sense [McC10, Hau06]. Services [Asp10]. Should [Sce18]. Smart [BSS18]. Software [Asp10, Ste18]. some [Sce18]. Something [Cas07]. Special [CPSB10]. Statistician [Mag06]. Statistics [Mag06]. Sterelny [Bro06a]. Structure [Mah10]. Structures [Mah10]. Sturgeon [Wei06]. System [Sce18].

Technology [Sha10]. Test [Pro06]. Theory [Cha06]. Thought [Ste03, Bro06a]. Tortoises [Bod07]. Totalisator [BC07, Dor07]. Totalisators [Dor18]. Transforming [Cha06]. Transistor [Asp07]. Tunny [Cop10]. Turing [Car10, CP12, Cop18, Dre10, Pro06, Ric06]. Type [Rou18]. Typewriter [CK06].

universal [Hau06]. University [Cla06]. Unlikely [BC07]. Unorganised [Rou18]. User [CK06]. User-friendly [CK06]. Using [Rou18].

Walter [Bod07]. world [Ste03, Bro06a]. wrong [Hau06].

Zealand [DT18, Tee07]. Zealand-Made [DT18].

References

Anderson:2005:DPH

[ADP06] David Anderson, Janet Delve, and Hans Pufal. Discovery of a prototype Hollerith machine in Paris. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380. URL http://

         rutherfordjournal.org/article010108.html.

Anonymous:2018:ARJ

[Ano18a] Anonymous. About the Rutherford Journal. Rutherford Journal, 5 (??):??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/editorial05.html.

Anonymous:2018:AC

[Ano18b] Anonymous. About the contributors. Rutherford Journal, 5(??): ??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/about05.html.

Aspray:2006:FJT

[Asp07] William Aspray. Den Fujita: From the Japanese transistor radio export business to Makudonaldo. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article020110.html.

Aspray:2010:RPR

[Asp10] William Aspray. The role of policy in the rise of the Indian software and IT-enabled services industry. Rutherford Journal, 3(??): ??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article030101.html.

Barrett:2005:JLP

[BC06] Lindsay Barrett and Matthew Connell. Jevons and the logic ‘piano’. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010103.html.

Barrett:2006:UHA

[BC07] Lindsay Barrett and Matthew Connell. An unlikely history of Australian computing: the reign of the totalisator. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL

http://rutherfordjournal.org/article020105.html.

Boden:2006:GWA

[Bod07] Margaret Boden. Grey Walter’s anticipatory tortoises. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article020101.html.

Bond:2018:FPO

[Bon18] Robin Bond. The figure of Prometheus and the origins of civilization. Rutherford Journal, 5(??):??, ???? 2018. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article050102.html.

Bowen:2010:BHE

[Bow10] Jonathan P. Bowen. A brief history of early museums online. Rutherford Journal, 3(??):??, ???? 2010. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article030103.html.

Browne:2005:EHM

[Bro06a] Derek Browne. The evolution of the human mind: Darwinian science or creation myth? Critical notice of Kim Sterelny, Thought in a Hostile World (2003). Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article010110.html.

Browne:2005:KLI

[Bro06b] Derek Browne. Konrad Lorenz on instinct and phylogenetic information. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010104.html.

Browne:2006:SMC

[Bro07] Derek Browne. A science of morality? Critical notice of Marc D. Hauser, Moral Minds (2006). Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article020107.html.

Bigelow:2018:OJJ

[BSS18] John Bigelow, Elizabeth Smart, and Helen Smart. Obituary of J. J. C. Smart, 1910–2012. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article050101.html.

Care:2006:CAC

[Car07] Charles Care. A chronology of analogue computing. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL

http://rutherfordjournal.org/article020106.html.

Carter:2010:TB

[Car10] Frank Carter. The Turing Bombe. Rutherford Journal, 3(??): ??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article030108.html.

Carpenter:2018:MMM

[Car18] Brian Carpenter. A meeting that missed its mark: The Paris Conference of 1951. Rutherford Journal, 5(??):??, ???? 2018. CO-

DEN ???? ISSN 1177-1380. URL http://rutherfordjournal.
org/article050103.html.

Casti:2006:WWS

[Cas07] John L. Casti. Why we see the events we do and not something else. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article020104.html.

Calude:2006:DMP

[CC07] Cristian S. Calude and Gregory J. Chaitin. A dialogue on mathematics and physics. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article020103.html.

Chalmers:2005:TAC

[Cha06] Alan Chalmers. Transforming atomic chemistry into an experimental science: The limitations of Dalton’s theory. Rutherford Journal, 1 (??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380. URL

http://rutherfordjournal.org/article010101.html.

Campbell-Kelly:2005:UFT

[CK06] Martin Campbell-Kelly. The user-friendly typewriter. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article010105.html.

Clarke:2005:RCU

[Cla06] Simon Clarke. Rutherford at Canterbury University College. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-

1380. URL http://rutherfordjournal.org/article010112.
html.

Copeland:2010:CBG

[Cop10] B. Jack Copeland. Colossus: Breaking the German ‘Tunny’ code at Bletchley Park. An illustrated history. Rutherford Journal, 3(??): ??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article030109.html.

Copeland:2018:ATL

[Cop18] Jack Copeland. Alan Turing’s lost notebook. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-

1380. URL http://www.math.utah.edu/pub/bibnet/authors/t/
turing-alan-mathison.bib;   http://rutherfordjournal.org/
article050104.html.

Corry:2010:HPN

[Cor10] Leo Corry. Hunting prime numbers — from human to electronic computers. Rutherford Journal, 3(??):??, ???? 2010. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article030105.html.

Copeland:2011:ATF

[CP12] B. Jack Copeland and Diane Proudfoot. Alan Turing and father of the modern computer. Rutherford Journal, 4(??):??, ???? 2011–2012. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article040101.html.

Copeland:2010:SIH

[CPSB10] B. Jack Copeland, Carl Posy, Oron Shagrir, and Parker Bright. A special issue on the history of computing. Rutherford Journal, 3(??): ??, ???? 2010. CODEN ???? ISSN 1177-1380.

Doran:2006:FAT

[Dor07] Bob Doran. The first automatic totalisator. Rutherford Journal, 2 (??):??, ???? 2006–2007. CODEN ???? ISSN 1177-1380. URL

http://rutherfordjournal.org/article020109.html.

Doran:2018:HHH

[Dor18] Bob Doran. Henry Hodsdon and his ’marble’ totalisators. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-1380. URL

http://rutherfordjournal.org/article050105.html.

Dresner:2010:TCM

[Dre10] Eli Dresner. Turing on computation, memory and behavior. Rutherford Journal, 3(??):??, ???? 2010. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article030104.html.

Doran:2018:POH

[DT18] Bob Doran and Andrew Trotman. Preserving our heritage: New Zealand-made computers. Rutherford Journal, 5(??):??, ???? 2018.

CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article050106.html.

Hauser:2006:MMH

[Hau06] Marc D. Hauser. Moral minds: how nature designed our universal sense of right and wrong. Ecco, New York, NY, USA, 2006. ISBN 0-06-078070-3. xx + 489 pp. LCCN BJ1012 .H348

2006. URL http://www.loc.gov/catdir/enhancements/fy0911/
2006041324-b.html;   http://www.loc.gov/catdir/enhancements/
fy0911/2006041324-d.html.

Harr:2005:DEF

[He06] Rom Harr and eacute. Do explanation formats in chemistry depend on agent causality? Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article010102.html.

Magnello:2005:KPO

[Mag06] M. Eileen Magnello. Karl Pearson and the origins of modern statistics: An elastician becomes a statistician. Rutherford Journal, 1(??): ??, ???? 2005–2006. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article010107.html.

Mahoney:2010:SCM

[Mah10] Michael S. Mahoney. The structures of computation and the mathematical structure of nature. Rutherford Journal, 3(??):??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article030107.html.

McCarthy:2010:FCS

[McC10] John McCarthy. Formalizing common sense knowledge in mathematical logic. Rutherford Journal, 3(??):??, ???? 2010. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article030110.html.

Musgrave:2006:LMA

[Mus07] Alan Musgrave. The ‘miracle argument’ for scientific realism. Rutherford Journal, 2(??):??, ???? 2006–2007. CODEN ???? ISSN 1177-

1380. URL http://rutherfordjournal.org/article020108.
html.

Numerico:2010:NML

[Num10] Teresa Numerico. The new machine: from logic to organization. Rutherford Journal, 3(??):??, ???? 2010. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article030102.html.

Proudfoot:2005:NIT

[Pro06] Diane Proudfoot. A new interpretation of the Turing Test. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article010113.html.

Richards:2005:TRM

[Ric06] Bernard Richards. Turing, Richards and morphogenesis. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article010109.html.

Rouly:2018:AIU

[Rou18] Ovi Chris Rouly. Artificial intelligence using P-type unorganised machines. Rutherford Journal, 5(??):??, ???? 2018. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article050107.html.

Scerri:2018:HSP

[Sce18] Eric Scerri. How should the periodic system be regarded? A brief look at some published proposals. Rutherford Journal, 5(??): ??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article050108.html.

Shamir:2010:CST

[Sha10] Eli Shamir. Computer science and technology in Israel, 1950– 1980. Rutherford Journal, 3(??):??, ???? 2010. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article030111.html.

Smart:2018:PMB

[Sma18] J. J. C. Smart. From physics to metaphysics and back again. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article050109.html.

Sterelny:2003:THW

[Ste03] Kim Sterelny. Thought in a hostile world: the evolution of human cognition. Blackwell, Malden, MA, USA, 2003. ISBN 0-631-18886-X, 0-631-18887-8 (paperback). xi + 262 pp. LCCN BF698.95 .S74 2003.

URL http://www.loc.gov/catdir/toc/fy044/2002155980.html.

Steinhart:2018:ES

[Ste18] Eric Steinhart. The existence of software. Rutherford Journal, 5 (??):??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article050110.html.

Swade:2010:ACC

[Swa10] Doron D. Swade. Automatic computation: Charles Babbage and computational method. Rutherford Journal, 3(??):??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://rutherfordjournal.

org/article030106.html.

Turing:2005:MCM

[T^{+}06] Alan Turing et al. The mind and the computing machine. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN ???? ISSN 1177-1380.

URL http://rutherfordjournal.org/article010111.html.

Tee:2006:SMP

[Tee07] Garry J. Tee. Science on the map: Places in New Zealand named after scientists. Rutherford Journal, 2(??):??, ???? 2006–2007. CO-

DEN ???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article020102.html.

Weierud:2005:BPS

[Wei06] Frode Weierud. Bletchley Park’s Sturgeon, the fish that laid no eggs. Rutherford Journal, 1(??):??, ???? 2005–2006. CODEN

???? ISSN 1177-1380. URL http://rutherfordjournal.org/
article010106.html.

Yock:2018:PPP

[Yoc18] Philip Yock. A perspective on particle physics. Rutherford Journal, 5(??):??, ???? 2018. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article050111.html.